OpenStudy (anonymous):

Solve the system of equations. y = 7x + 43 y = 2x + 13 a. ( -6, 1) c. ( 6, 1) b. ( 6, -1) d. No solution

OpenStudy (anonymous):

@amistre64 @AravindG @agent0smith @arabpride @aprilann143 @AuroraB @bobsue @baby456 @bossman101 @Blank  @BMasterOB123 @BMasterOB123 @bobsue @BasedGod1122 @g.u.y. @geo12345679789

OpenStudy (anonymous):

@HelloGoodmorning @helpout3 @HMT

OpenStudy (aravindg):

Subtract both equations.

OpenStudy (anonymous):

so 5 and then 30

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

Take 30 to the other side 5x=-30 x=?

OpenStudy (anonymous):

-6

OpenStudy (aravindg):

Yep :)

OpenStudy (anonymous):

y = -5x + 24 y = 4x - 21 a. ( -5, -1) c. ( -1, 5) b. ( 5, -1) d. No solution

OpenStudy (anonymous):

can you help me work through this one?

OpenStudy (aravindg):

y = 2x + 13 Put x=-6

OpenStudy (aravindg):

First lets finish this question.

OpenStudy (blank ):

Do you want to use the substitution method or the Addition/Elimination method?

OpenStudy (anonymous):

ist (-6,1)

OpenStudy (aravindg):

Yep :)

OpenStudy (blank ):

Yep, it's A. :)

OpenStudy (anonymous):

y = -6x + 9 y = -3x + 3

OpenStudy (aravindg):

subtract both as before.

OpenStudy (anonymous):

-3 and 6

OpenStudy (aravindg):

-3x+6=0 x=?

OpenStudy (anonymous):

(2,-3)

OpenStudy (aravindg):

Good work!!!

OpenStudy (anonymous):

The sum of two numbers is 98. Their difference is 22. Write a system of equations that describes this situation. Solve by elimination to find the two numbers. a. x + y = 98 x – y = 22 55 and 33 c. x – y = 98 x + y = 22 59 and 39 b. x + y = 98 x – y = 22 60 and 38 d. x + y = 22 y – x = 98 55 and 39

OpenStudy (aravindg):

Please post your questions separately. Don't use a single thread for all your questions. Close this question and post this separately.

OpenStudy (texaschic101):

x + y = 98 x - y = 22 -------------add 2x = 120 x = 60 now sub 60 in for x in either of the equations to find y can you do that ?

OpenStudy (blank ):

I did the work already and I got b. Your b and c are messed up a little by the way. XD x=60 y=38

OpenStudy (texaschic101):

lol...I noticed that too...she went from A to C to B to D